2x^ 3 y^ prime =y(y^ 2 +3x^ 2 )



1
Expert's answer
2021-12-07T11:39:13-0500

"2x^ 3 y' =y(y^ 2 +3x^ 2 )"

"y=tx"

"y'=t'x+t"


"2x^ 3 (t'x+t) =tx((tx)^ 2 +3x^ 2 )"

"2x^4t'+2x^3t=t^3x^3+3tx^3"

"2xt'+2t=t^3+3t"


"\frac{dt}{t^3+t}=\frac{dx}{2x}"


"-\frac{ln(1/t^2+1)}{2}=\frac{lnx}{2}+lnc/2"


"\frac{1}{1/t^2+1}=cx"


"\frac{y^2}{x^2+y^2}=cx"